# The ECME Algorithm Using Factor Analysis for DOA Estimation in Nonuniform Noise

### v1

@@ IV-A2 E-step
where $\mathbf{C}^{(k)}=\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}+\mathbf{Q}^{(k-1)}$, $\boldsymbol{\delta}^{(k)}=[\mathbf{C}^{(k)}]^{-1}\mathbf{Q}^{(k-1)}$, and $\boldsymbol{\Delta}^{(k)}=\mathbf{Q}^{(k-1)}-\mathbf{Q}^{(k-1)}\boldsymbol{\delta}^{(k)}\succeq\mathbf{0}_{N}$ [^4] We do not adopt $[\mathbf{C}^{(k)}]^{-1}=\mathbf{q}^{(k)}-\mathbf{q}^{(k)}\mathbf{S}^{(k)}(\boldsymbol{\Lambda}^{(k)}+\mathbf{I}_{M})^{-1}[\mathbf{S}^{(k)}]^{H}\mathbf{q}^{(k)}$ with $\mathbf{q}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1}$ since numerical results show its low accuracy.. The conditional distribution of $\mathbf{x}(t)$ can be derived from [16], and
@@
IV-A3 Second CM-step
This step only obtains $\mathbf{Q}^{(k)}$ by maximizing $l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ in (16) with respect to $\mathbf{Q}$, i.e.,
@@ IV-A3 Second CM-step
1: Initialize $\mathbf{Q}^{(0)}$, $K$, and $k=1$.
@@
4:  Obtain $\mathbf{Q}^{(k)}$ using (IV-A3).
@@
IV-B Convergence and Complexity
From [5, 17], we know that the ECME algorithm satisfies certain regularity conditions and its sequence $\{(\mathbf{S}^{(k)},\mathbf{Q}^{(k)})\}$ stably converges to a stationary point of $f(\mathbf{S},\mathbf{Q})$. Moreover, Eq. (15) and Eq. (IV-A3) guarantee that the solution provided by the ECME algorithm is feasible for problem (5).
By comparing Algorithm 1 and Algorithm 2, we can find that at the $k$th iteration, the computational complexity of Algorithm 1 lies in the eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}\in\mathbb{C}^{N\times N}$ in (9) and the inverse of $\big{(}\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}+\mathbf{I}_{N}\big{)}\in\mathbb{C}^{N\times N}$ in (11). Also, the computational complexity of Algorithm 2 lies in the eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}\in\mathbb{C}^{N\times N}$ in (15) and the inverse of $\mathbf{C}^{(k)}\in\mathbb{C}^{N\times N}$ in (IV-A3). Thus, the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method.
V Numerical Results
@@
After obtaining $\hat{\mathbf{S}}$ or the signal subspace estimate $\mathcal{C}(\hat{\mathbf{S}})$, we perform the eigen-decomposition of $\hat{\mathbf{S}}\hat{\mathbf{S}}^{H}$ to obtain $\hat{\mathbf{U}}_{n}\in\mathbb{C}^{N\times(N-M)}$, which consists of the eigenvectors corresponding to the $N-M$ smallest eigenvalues. To proceed, we adopt the orthogonal projector $\hat{\boldsymbol{\Pi}}_{n}=\hat{\mathbf{U}}_{n}\hat{\mathbf{U}}_{n}^{H}$ [^5] We do not adopt $\hat{\boldsymbol{\Pi}}_{n}=\mathbf{I}_{N}-\hat{\mathbf{S}}(\hat{\mathbf{S}}^{H}\hat{\mathbf{S}})^{-1}\hat{\mathbf{S}}^{H}$ since numerical results show its low accuracy. onto the noise subspace estimate $\mathcal{C}(\hat{\mathbf{U}}_{n})$ and the Root-MUSIC algorithm [18] to obtain the DOA estimates $\hat{\theta}_{m}$’s.
Fig. 2 shows a scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from both algorithms under 100100100 independent realizations. Here, $L=100$ and $K=K_{1}=1000$. The same snapshots of each realization are processed by both algorithms. We can know that under the same number of iterations $K=K_{1}$, both algorithms provide different $\hat{\mathbf{S}}$’s and for each realization, the $\hat{\mathbf{S}}\hat{\mathbf{S}}^{H}$ provided by the ECME algorithm is closer to the solution of $\mathbf{S}\mathbf{S}^{H}$ in (5). Hence, Fig. 3 compares the root mean square error (RMSE) performances of DOA estimation related to both algorithms with the Cramer–Rao lower bound (CRLB) [19, 20]. Here, $K=K_{1}=100$ and each RMSE is based on 100010001000 independent realizations. We observe that the RMSEs of DOA estimation from both algorithms approach the CRLBs as $L$ increases. More importantly, since the ECME algorithm yields faster convergence, its $\mathcal{C}(\mathcal{\hat{\mathbf{S}}})$ is more accurate under $K=K_{1}$ and hence its RMSE performance of DOA estimation is better.
@@
VI Conclusion
In this letter, the ECME algorithm, an extension of the EM algorithm, has been designed for MLFA, which has almost the same computational complexity at each iteration as the FAAN method with excellent convergence properties. However, numerical results have shown that the ECME algorithm yields faster stable convergence and is computationally more efficient. Importantly, the signal subspace estimated by the ECME algorithm can be used for the subspace based DOA estimation in unknown nonuniform noise.

### v2

@@ 4.1.2 E-step:
where $\mathbf{C}^{(k)}=\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}+\mathbf{Q}^{(k-1)}$, $\boldsymbol{\delta}^{(k)}=[\mathbf{C}^{(k)}]^{-1}\mathbf{Q}^{(k-1)}$ [^4] We do not adopt $\mathbf{C}^{(k)}]^{-1}=\mathbf{q}^{(k)}-\mathbf{q}^{(k)}\mathbf{S}^{(k)}(\boldsymbol{\Lambda}^{(k)}+\mathbf{I}_{M})^{-1}[\mathbf{S}^{(k)}]^{H}\mathbf{q}^{(k)}$ with $\mathbf{q}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1}$ since numerical results show its low accuracy., and $\boldsymbol{\Delta}^{(k)}=\mathbf{Q}^{(k-1)}-\mathbf{Q}^{(k-1)}\boldsymbol{\delta}^{(k)}\succeq\mathbf{0}_{N}$. The conditional distribution of $\mathbf{v}(t)$ can be derived from [20], and
@@
4.1.3 Second CM-step:
This step only obtains $\mathbf{Q}^{(k)}$ by maximizing $l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ in (4.1.2) with respect to $\mathbf{Q}$, i.e.,
@@ 4.1.3 Second CM-step:
1: Initialize $\mathbf{Q}^{(0)}\succ\mathbf{0}_{N}$, $K$, and $k=1$.
@@
4:  Obtain $\mathbf{Q}^{(k)}$ using (4.1.3).
@@
4.2 Convergence and Complexity
From [7, 21], we know that the ECME algorithm satisfies certain regularity conditions and its sequence $\big{\{}\big{(}\mathbf{S}^{(k)},\mathbf{Q}^{(k)}\big{)}\big{\}}$ always converges to a stationary point of $f(\mathbf{S},\mathbf{Q})$. Fortunately, simulation results in [10] and the subsequent section show that although the objective function $f(\mathbf{S},\mathbf{Q)}$ in problem (5) is non-convex, it seems to have a favourable landscape, which makes its minimization by the FAAN method or the ECME algorithm a relatively straightforward task. Moreover, Eq. (15) and Eq. (4.1.3) guarantee that the solution provided by the ECME algorithm is “proper" for problem (5) [18, 22].
By comparing Algorithm 1 and Algorithm 2, we can find that at the $k$th iteration, the computational complexity of Algorithm 1 lies in the eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}\in\mathbb{C}^{N\times N}$ in (9) and the inverse of $\big{(}\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}+\mathbf{I}_{N}\big{)}\in\mathbb{C}^{N\times N}$ in (3.2). Also, the computational complexity of Algorithm 2 lies in the eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}\in\mathbb{C}^{N\times N}$ in (15) and the inverse of $\mathbf{C}^{(k)}\in\mathbb{C}^{N\times N}$ in (4.1.3). Thus, the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method.
5 Numerical Results
@@
[…]
@@
6 Conclusion
In this article, the ECME algorithm, an extension of the EM algorithm, has been designed for MLFA, which has almost the same computational complexity at each iteration as the FAAN method with excellent convergence properties. However, numerical results have shown that the ECME algorithm yields faster stable convergence and is computationally more efficient. Importantly, the signal subspace estimated by the ECME algorithm can be employed for the subspace based DOA estimation in unknown nonuniform noise.